# Département de mathématiques Université du Québec à Montréal 201, Avenue du Président-Kennedy

### v2

@@ 4.2 Empirical power under $\mathcal{H}_{1}$
The study is conducted with a sample size of $n=50$ and a significance level of $\alpha=0.05$. We evaluate empirical powers across three families of null distributions: (1) $\mathcal{H}_{0}:X_{i}\sim\mathrm{normal}(\mu_{0},\sigma_{0})$, (2) $\mathcal{H}_{0}:X_{i}\sim\mathrm{Student^{\prime}s}\,t(\lambda_{0}=4,\mu_{0},\sigma_{0})$, and (3) $\mathcal{H}_{0}:X_{i}\sim\mathrm{exponential}(\beta_{0})$, with all parameters treated as unknown except for the Student’s $t$ distribution, where $\lambda_{0}=4$ is assumed known. The unknown nuisance parameters are replaced by their maximum likelihood estimators under the null hypothesis; see Table 2. Critical values for each of the three families were computed using 1,000,000 simulations for each of the six goodness-of-fit tests, for $n=50$ and $\alpha=0.05$. They are reported in Table 6. We observe that the simulated critical values for the LK test are very close to the chi-square quantile $\chi^{2}_{2,\,0.95}=5.99146$, and those of the $T_{n}$ test are even closer. This again illustrates that the chi-square approximation is excellent for both $T_{n}$ and LK, even for relatively small samples such as $n=50$.
@@
We then summarized the overall behavior by averaging the resulting powers, i.e., by computing the area under each power curve divided by the length of the parameter range, and these averages are reported in Tables 78 and 9. The full power curves themselves are provided in Section C.15 of Appendix C. The overall summary of empirical powers across the three families of null distributions is presented in Table 10. For each test, the average power over the ten families of alternatives was computed to obtain a single overall measure. The AD test achieves the highest mean power at 57.4%, followed closely by the $T_{n}$ test at 57.2% (a gap of only 0.2%), with LK in third place at 54.7% (a gap of 2.7% relative to AD). As mentioned above, these results illustrate that the $T_{n}$ and LK tests perform well, while recognizing that definitive conclusions would require a more extensive study.
@@
$\mathrm{EPD}(0.3\leq\lambda\leq 1.2,\mu,\sigma)$ 43.0 40.3 40.6 41.8 39.4 41.7
$\mathrm{EPD}(1.2\leq\lambda\leq 15,\mu,\sigma)$ 62.2 55.5 51.3 49.2 40.9 51.1
$\mathrm{SN}(0\leq\lambda\leq 16,\mu,\sigma)$ 49.2 49.4 56.6 45.3 37.1 43.1
Average power 51.5 48.4 49.5 45.4 39.1 45.3
Gap with best 0.0 3.1 2.0 6.0 12.3 6.1
Table 8: Average empirical powers (%) for testing $\mathcal{H}_{0}:X_{i}\sim\mathrm{Student^{\prime}s}\,t(\lambda_{0}=4\,\text{known},\mu_{0},\sigma_{0}\ \text{unknown})$ under EPD and skew normal (SN) alternatives, for a sample size of $n=50$ and a significance level of $0.05$. The power results are invariant to the specific values of $\mu$ and $\sigma$ under $\mathcal{H}_{1}$.
@@
Overall average power 57.2 54.7 57.4 53.1 50.0 54.0
Gap with best 0.2 2.7 0.0 4.3 7.4 3.4
@@
When examining the tests individually, the two highest average powers for normality are obtained by AD (42.4%) and $T_{n}$ (40.7%). For the Student’s $t_{4}$ test, the best performances are achieved by $T_{n}$ (51.5%) and AD (49.5%). For the exponential test, AD (74.6%) and $T_{n}$ (73.8%) again yield the strongest results. Interestingly, the gain in power obtained by using the $T_{n}$ test over the LK test averages 2.5% across all settings, with respective improvements of 1.4%, 3.1%, and 2.9% for the normality, Student’s $t_{4}$, and exponential tests.
@@ 4.3 Empirical power for the Laplace distribution
In this section, we revisit the comprehensive simulation study of goodness-of-fit tests for the Laplace distribution presented by Desgagné et al. [9], assuming that both $\mu$ and $\sigma$ are unknown. This Monte Carlo study evaluated $40$ competing goodness-of-fit tests against a broad set of $400$ alternative distributions, including both symmetric and asymmetric cases with light and heavy tails, for sample sizes $n\in\{20,50,100,200\}$ and a significance level $\alpha=0.05$. In particular, it included the MM-based LK test proposed by Langholz and Kronmal [17].
Our new MM-based Laplace test $T_{n}$ was added to this set of contenders, increasing the total number of tests considered to $41$. We deliberately chose the MM-based version of $T_{n}$, rather than its ML-based counterpart, in order to ensure consistency with the MM-based LK test already included in the original study.
Table 11 summarizes the results, listing the $12$ most powerful tests for each sample size. To aid interpretation, the table reports the “gap,” defined as the difference between a given test’s average power and the highest average power achieved among all $41$ tests for the corresponding sample size. The results clearly demonstrate the strength of our new procedure. Our test $T_{n}$ (highlighted in dark gray) ranks 12th, 4th, 1st, and 1st out of 41 contenders for the sample sizes $n=20$, 50, 100, and 200, respectively. The last two rows of Table 11 summarize the gaps across the four sample sizes considered for each of the leading tests, reporting both the maximum and the average of these four gaps. Overall, when all sample sizes are considered, our test $T_{n}$ emerges as the most powerful test on average among all 41 competitors, with an average power gap of $1.2\%$. Furthermore, $T_{n}$ improves average power by margins of 1.5%, 3.2%, 3.1%, and 1.8% compared to the LK test (in light gray), which ranks 16th, 9th, 3rd, and 4th, respectively, for $n=20$, 50, 100, and 200.
@@ 5 Asymptotic power under sequences of local alternatives
This section examines the asymptotic power of the test statistic $T_{n}(\hat{\bm{\theta}}_{n})$ under sequences of local alternatives. Asymptotic power curves are provided for the gamma and exponential power distribution (EPD) tests with unknown nuisance parameters. In each case, a sequence of local alternatives is constructed by embedding the null distribution within a more general parametric family, which allows us to apply the theoretical framework of Desgagné et al. [11, Theorem 2] and derive the noncentrality parameter of the exact limiting chi-square distribution of the test statistic. The performance of the test $T_{n}(\hat{\bm{\theta}}_{n})$ is compared with that of a natural benchmark provided by Rao’s score test (also known as the Lagrange multiplier test), as well as with its asymptotically equivalent counterpart, the generalized likelihood ratio test (GLRT), for which the exact limiting distribution is also derived under the null hypothesis and under sequences of local alternatives.
@@
Under $\mathcal{H}_{0}$ and as $n\to\infty$, we obtain
$\mathcal{R}_{n},\Lambda_{n}\rightsquigarrow\chi_{\dim(\bm{\phi}_{0})}^{2}.$
Applying now the theoretical framework of Desgagné et al. [11, Theorem 2], we derive the asymptotic distribution of the tests $T_{n}(\hat{\bm{\theta}}_{n})$, $\mathcal{R}_{n}$ and $\Lambda_{n}$ under sequences of local alternatives.
For the ML-based $T_{n}(\hat{\bm{\theta}}_{n})$, consider the matrix $M(\bm{\phi}_{0},\bm{\theta}_{0})$ defined by
@@
The vector $\bm{\tau}$ denotes the kernel defined in (2), and $G(\bm{\theta}_{0})$ is the score–kernel cross-moment matrix defined in (3).
Similarly for the ML-based tests $\mathcal{R}_{n}$ and $\Lambda_{n}$, by using instead the kernel $\bm{\tau}_{\mathcal{R}}(x,\bm{\phi}_{0},\bm{\theta}_{0})$, we obtain
@@
Note that for the maximum likelihood case, that is, $\hat{\bm{\theta}}_{n}=\tilde{\bm{\theta}}_{n}$, we have $\bm{r}(x,\bm{\theta}_{0})=\bm{s}(x,\bm{\theta}_{0})\equiv\bm{s}_{\theta}(x,\bm{\phi}_{0},\bm{\theta}_{0})$. In this case, $R(\bm{\theta}_{0})=I(\bm{\theta}_{0})\equiv I_{\theta,\theta}(\bm{\phi}_{0},\bm{\theta}_{0})$, $S_{\phi,\theta}(\bm{\phi}_{0},\bm{\theta}_{0})=I_{\phi,\theta}(\bm{\phi}_{0},\bm{\theta}_{0})$ and $M(\bm{\phi}_{0},\bm{\theta}_{0})$ defined in (10) is equivalent to the form in (9).
By applying Theorem 2 of Desgagné et al. [11], we obtain, under $\mathcal{H}_{1,n}(\bm{\delta})$ and as $n\to\infty$,
@@
Within this framework, the quantities required to compute $V(\bm{\phi}_{0},\bm{\theta}_{0})$ and $V_{\mathcal{R}}(\bm{\phi}_{0},\bm{\theta}_{0})=\Sigma_{\mathcal{R}}(\bm{\phi}_{0},\bm{\theta}_{0})$, and ultimately the noncentrality parameters, are derived for the gamma and EPD tests under specific sequences of local alternatives in order to study the asymptotic power of the test statistic $T_{n}(\hat{\bm{\theta}}_{n})$ and of the benchmark tests $\Lambda_{n}$ and $\mathcal{R}_{n}$.
@@
See Section C.13 of Appendix C.
@@
where $\bm{\delta}=[\delta_{1},\delta_{2}]^{\top}\in\mathbb{R}^{2}\setminus\{[0,0]^{\top}\}$ is fixed. The fixed parameter vector is $\phi_{0}=[\alpha_{0},\rho_{0}]^{\top}=[1/2,\lambda_{0}]^{\top}$, while the unknown parameter vector is $\bm{\theta}_{0}=[\mu_{0},\sigma_{0}]^{\top}$. Consider the ML-based test $T_{n}(\hat{\mu}_{n},\hat{\sigma}_{n})$, Rao’s score test $\mathcal{R}_{n}$, and the GLRT $\Lambda_{n}$. Then, we have
$G(\bm{\theta}_{0})=\begin{bmatrix}0~&~h_{1}(\lambda_{0})\\[2.84526pt] \frac{h_{2}(\lambda_{0})}{\lambda_{0}^{1/\lambda_{0}-1}\Gamma(1/\lambda_{0})}~&~0\end{bmatrix},\qquad G_{\bm{\phi}}(\bm{\phi_{0}},\bm{\theta}_{0})=\begin{bmatrix}0&-\frac{1}{\lambda_{0}^{2}}h_{3}(\lambda_{0})\\[2.84526pt] -2h_{35}(\lambda_{0})&0\end{bmatrix}$,
$I(\bm{\theta}_{0})=\begin{bmatrix}\frac{\lambda_{0}^{2-2/\lambda_{0}}\Gamma(2-1/\lambda_{0})}{\Gamma(1/\lambda_{0})}~&~0\\[2.84526pt] 0~&~\lambda_{0}\end{bmatrix},\quad I_{\phi,\phi}(\bm{\phi}_{0},\bm{\theta}_{0})=\begin{bmatrix}4(\lambda_{0}+1)~&~0\\[2.84526pt] 0~&~\frac{1}{\lambda_{0}^{3}}\left\{\Big(\frac{1}{\lambda_{0}}+1\Big)\psi_{1}\Big(\frac{1}{\lambda_{0}}+1\Big)-1+(C_{\lambda_{0}}+1)^{2}\right\}\end{bmatrix}$,
$I_{\phi,\theta}(\bm{\phi}_{0},\bm{\theta}_{0})=\begin{bmatrix}\frac{-2\lambda_{0}^{2-1/\lambda_{0}}}{\Gamma(1/\lambda_{0})}&0\\[2.84526pt] 0&-\frac{1}{\lambda_{0}}(C_{\lambda_{0}}+1)\end{bmatrix},\quad\Sigma_{\mathcal{R}}(\bm{\phi}_{0},\bm{\theta}_{0})=\begin{bmatrix}4(\lambda_{0}+1)-\displaystyle\frac{4\lambda_{0}^{2}}{\Gamma(2-1/\lambda_{0})\,\Gamma(1/\lambda_{0})}&0\\[2.84526pt] 0&\frac{\big(\frac{1}{\lambda_{0}}+1\big)\psi_{1}\big(\tfrac{1}{\lambda_{0}}+1\big)-1}{\lambda_{0}^{3}}\end{bmatrix}$,
where $C_{\lambda_{0}}=\psi(1/\lambda_{0}+1)+\ln(\lambda_{0})$, $\Gamma(\cdot)$, $\psi(\cdot)$ and $\psi_{1}(\cdot)$ denote the gamma, digamma and trigamma functions, respectively, and the functions $h_{i}(\cdot)$ are defined in Table 13.
Consider the MM-based test $T_{n}(\hat{\mu}_{n},\hat{\sigma}_{n})$. Then, we have
$R(\bm{\theta}_{0})=\begin{bmatrix}\frac{\Gamma(1/\lambda_{0})}{\lambda_{0}^{2/\lambda_{0}}\Gamma(3/\lambda_{0})}&0\\[2.84526pt] 0&4D_{\lambda_{0}}\end{bmatrix},\qquad S_{\phi,\theta}(\bm{\phi}_{0},\bm{\theta}_{0})=\begin{bmatrix}\frac{-4\Gamma(2/\lambda_{0})}{\lambda_{0}^{1/\lambda_{0}}\Gamma(3/\lambda_{0})}&0\\[2.84526pt] 0&-\frac{2D_{\lambda_{0}}}{\lambda_{0}^{2}}\left\{2\ln(\lambda_{0})+3\psi(3/\lambda_{0})-\psi(1/\lambda_{0})\right\}\end{bmatrix}$,
where $D_{\lambda_{0}}$ is defined in the second row of Table 3. Note that the resulting matrices $V(\bm{\phi}_{0},\bm{\theta}_{0})$ (for both the ML and MM cases) and $\Sigma_{\mathcal{R}}(\bm{\phi}_{0},\bm{\theta}_{0})$ depend only on the value of $\lambda_{0}$. Therefore, to simplify the expressions, we set $\mu_{0}=0,\sigma_{0}=1$.
@@
See Section C.14 of Appendix C.
@@
Note that the Rao’s score test for the EPD model coincides with the goodness-of-fit test based on the $\lambda$-th power skewness and kurtosis introduced by Desgagné et al. [10]. Since the family of local alternatives is the APD, which accommodates a wide range of asymmetry and tail thickness, this test can be used as an omnibus procedure even though it is specifically designed for this APD alternative. This helps explain why the power curves of the $T_{n}$ tests and of the GLRT/score-test benchmark are much closer for the EPD case than for the gamma case, for which the family of alternatives is more restrictive.
6 Real-data application

### v3

@@ 4.2 Empirical power under $\mathcal{H}_{1}$
The study is conducted with a sample size of $n=50$ and a significance level of $\alpha=0.05$. We evaluate empirical powers across three families of null distributions: (1) $\mathcal{H}_{0}:X_{i}\sim\mathrm{normal}(\mu_{0},\sigma_{0})$, (2) $\mathcal{H}_{0}:X_{i}\sim\mathrm{Student^{\prime}s}\,t(\lambda_{0}=2,\mu_{0},\sigma_{0})$, and (3) $\mathcal{H}_{0}:X_{i}\sim\mathrm{exponential}(\beta_{0})$, with all parameters treated as unknown except for the Student’s $t$ distribution, where $\lambda_{0}=2$ is assumed known. The unknown nuisance parameters are replaced by their maximum likelihood estimators under the null hypothesis; see Table 12. Critical values for each of the three families were computed using 1,000,000 simulations for each of the six goodness-of-fit tests, for $n=50$ and $\alpha=0.05$. They are reported in Table 5. We observe that the simulated critical values for the LK test are very close to the chi-square quantile $\chi^{2}_{2,\,0.95}=5.99146$, and those of the $T_{n}$ test are even closer. This again illustrates that the chi-square approximation is excellent for both $T_{n}$ and LK, even for relatively small samples such as $n=50$.
@@
We then summarized the overall behavior by averaging the resulting powers, i.e., by computing the area under each power curve divided by the length of the parameter range, and these averages are reported in Tables 67 and 8. The full power curves themselves are provided in Section S15 of the Omnibus goodness-of-fit tests for univariate continuous distributions based on trigonometric moments. To aid interpretation, the tables report the “gap with best,” defined as the difference between a given test’s average power and the highest average power achieved among all $6$ tests. The overall summary of empirical powers across the three families of null distributions is presented in Table 9. For each test, the average power over the ten families of alternatives was computed to obtain a single overall measure. The $T_{n}$ test achieves the highest mean power at 59.9%, followed by three close competitors: the LK test at 56.9% (a gap of 3.0%), the Watson test at 56.6% (a gap of 3.3%), and the AD test at 55.9% (a gap of 4.0%). As mentioned above, these results illustrate that the $T_{n}$ and LK tests perform well, while recognizing that definitive conclusions would require a more extensive study.
When examining the tests individually, the two highest average powers for normality are obtained by AD (42.4%) and $T_{n}$ (40.7%). For the Student’s $t_{2}$ test, the best performance is achieved by $T_{n}$ (60.6%), followed by LK (55.8%) and Watson (54.0%). For the exponential test, AD (74.6%) and $T_{n}$ (73.8%) again yield the strongest results. Interestingly, the gain in power obtained by using the $T_{n}$ test over the LK test averages 3.0% across all settings, with respective improvements of 1.4%, 4.8%, and 2.9% for the normality, Student’s $t_{2}$, and exponential tests.
@@
$\mathrm{EPD}(0.25\leq\lambda\leq 0.8,\mu,\sigma)$ 44.5 40.2 34.0 37.7 41.3 45.1
$\mathrm{EPD}(1.2\leq\lambda\leq 10,\mu,\sigma)$ 72.3 65.2 43.1 42.1 49.3 62.7
$\mathrm{SN}(0\leq\lambda\leq 16,\mu,\sigma)$ 64.9 62.0 56.7 55.8 45.7 54.1
Average power 60.6 55.8 44.6 45.2 45.5 54.0
Gap with best 0.0 4.8 16.0 15.4 15.1 6.6
Table 7: Average empirical powers (%) for testing $\mathcal{H}_{0}:X_{i}\sim\mathrm{Student^{\prime}s}\,t(\lambda_{0}=2\,\text{known},\mu_{0},\sigma_{0}\ \text{unknown})$ under EPD and skew normal (SN) alternatives, for a sample size of $n=50$ and a significance level of $0.05$. The power results are invariant to the specific values of $\mu$ and $\sigma$ under $\mathcal{H}_{1}$.
@@
Overall average power 59.9 56.9 55.9 53.0 51.9 56.6
Gap with best 0.0 3.0 4.0 6.9 8.0 3.3
@@ 4.3 Empirical power for the Laplace distribution
In this section, we revisit the comprehensive simulation study of goodness-of-fit tests for the Laplace distribution presented by Desgagné et al. [9], assuming that both $\mu$ and $\sigma$ are unknown. This Monte Carlo study evaluated $40$ competing goodness-of-fit tests against a broad set of $400$ alternative distributions, including both symmetric and asymmetric cases with light and heavy tails, for sample sizes $n\in\{20,50,100,200\}$ and a significance level $\alpha=0.05$. In particular, it included the MM-based LK test proposed by Langholz and Kronmal [17]. Our new MM-based Laplace test $T_{n}$ was added to this set of contenders, increasing the total number of tests considered to $41$. We deliberately chose the MM-based version of $T_{n}$, rather than its ML-based counterpart, in order to ensure consistency with the MM-based LK test already included in the original study.
Table 10 summarizes the results, listing the $12$ most powerful tests for each sample size. The results clearly demonstrate the strength of our new procedure. Our test $T_{n}$ (highlighted in dark gray) ranks 12th, 4th, 1st, and 1st out of 41 contenders for the sample sizes $n=20$, 50, 100, and 200, respectively. The last two rows of Table 10 summarize the gaps across the four sample sizes considered for each of the leading tests, reporting both the maximum and the average of these four gaps. Overall, when all sample sizes are considered, our test $T_{n}$ emerges as the most powerful test on average among all 41 competitors, with an average power gap of $1.2\%$. Furthermore, $T_{n}$ improves average power by margins of 1.5%, 3.2%, 3.1%, and 1.8% compared to the LK test (in light gray), which ranks 16th, 9th, 3rd, and 4th, respectively, for $n=20$, 50, 100, and 200.
@@ 5 Asymptotic power under sequences of local alternatives
This section examines the asymptotic power of the test statistic $T_{n}(\hat{\bm{\theta}}_{n})$ under sequences of local alternatives. Asymptotic power curves are provided for the gamma and exponential power distribution (EPD) tests with unknown nuisance parameters. In each case, a sequence of local alternatives is constructed by embedding the null distribution within a more general parametric family, which allows us to apply the theoretical framework of Desgagné et al. [11, Theorem 2] and derive the noncentrality parameter of the limiting chi-square distribution of the test statistic. The performance of the test $T_{n}(\hat{\bm{\theta}}_{n})$ is compared with that of a natural benchmark provided by Rao’s score test (also known as the Lagrange multiplier test), as well as with its asymptotically equivalent counterpart, the generalized likelihood ratio test (GLRT), for which the limiting distribution is also derived under the null hypothesis and under sequences of local alternatives.
@@
Under $\mathcal{H}_{0}$ and as $n\to\infty$, we obtain $\mathcal{R}_{n},\Lambda_{n}\rightsquigarrow\chi_{\dim(\bm{\phi}_{0})}^{2}.$
Applying now the theoretical framework of Desgagné et al. [11, Theorem 2], we derive the asymptotic distribution of the tests $T_{n}(\hat{\bm{\theta}}_{n})$, $\mathcal{R}_{n}$ and $\Lambda_{n}$ under sequences of local alternatives. For the ML-based $T_{n}(\hat{\bm{\theta}}_{n})$, consider the matrix $M(\bm{\phi}_{0},\bm{\theta}_{0})$ defined by
@@
The vector $\bm{\tau}$ denotes the kernel defined in (2), and $G(\bm{\theta}_{0})$ is the score–kernel cross-moment matrix defined in (3). Similarly for the ML-based tests $\mathcal{R}_{n}$ and $\Lambda_{n}$, by using instead the kernel $\bm{\tau}_{\mathcal{R}}(x,\bm{\phi}_{0},\bm{\theta}_{0})$, we obtain
@@
Note that for the maximum likelihood case, that is, $\hat{\bm{\theta}}_{n}=\tilde{\bm{\theta}}_{n}$, we have $\bm{r}(x,\bm{\theta}_{0})=\bm{s}(x,\bm{\theta}_{0})\equiv\bm{s}_{\theta}(x,\bm{\phi}_{0},\bm{\theta}_{0})$. In this case, $R(\bm{\theta}_{0})=I(\bm{\theta}_{0})\equiv I_{\theta,\theta}(\bm{\phi}_{0},\bm{\theta}_{0})$, $S_{\phi,\theta}(\bm{\phi}_{0},\bm{\theta}_{0})=I_{\phi,\theta}(\bm{\phi}_{0},\bm{\theta}_{0})$ and $M(\bm{\phi}_{0},\bm{\theta}_{0})$ defined in (10) is equivalent to the form in (9). By applying Theorem 2 of Desgagné et al. [11], we obtain, under $\mathcal{H}_{1,n}(\bm{\delta})$ and as $n\to\infty$,
@@
Within this framework, the quantities required to compute $V(\bm{\phi}_{0},\bm{\theta}_{0})$ and $V_{\mathcal{R}}(\bm{\phi}_{0},\bm{\theta}_{0})=\Sigma_{\mathcal{R}}(\bm{\phi}_{0},\bm{\theta}_{0})$, and ultimately the noncentrality parameters, are derived for the gamma and EPD tests under specific sequences of local alternatives in order to study the asymptotic power of the test statistic $T_{n}(\hat{\bm{\theta}}_{n})$ and the benchmark tests $\Lambda_{n}$ and $\mathcal{R}_{n}$.
@@
This essentially consists of decomposing the matrices $G(\lambda_{0},1,1)$ and $I(\lambda_{0},1,1)$ of the GG distribution from Table 2 according to the fixed and unknown parameters; see Section S13 of the Omnibus goodness-of-fit tests for univariate continuous distributions based on trigonometric moments for full details.
@@
where $\bm{\delta}=[\delta_{1},\delta_{2}]^{\top}\in\mathbb{R}^{2}\setminus\{[0,0]^{\top}\}$ is fixed. The fixed parameter vector is $\phi_{0}=[\alpha_{0},\rho_{0}]^{\top}=[1/2,\lambda_{0}]^{\top}$, while the unknown parameter vector is $\bm{\theta}_{0}=[\mu_{0},\sigma_{0}]^{\top}$. Consider the ML-based and MM-based tests $T_{n}(\hat{\mu}_{n},\hat{\sigma}_{n})$, Rao’s score test $\mathcal{R}_{n}$, and the GLRT $\Lambda_{n}$. Then, we have $G(\bm{\theta}_{0})=[G(\lambda_{0},0,1)]_{1:2,2:3}$, $I(\bm{\theta}_{0})=[I(\lambda_{0},0,1)]_{2:3,2:3}$, and $R(\bm{\theta}_{0})=R_{\lambda_{0}}(0,1)$, using the matrices of the EPD distribution (ML and MM cases) from Table 2, and
$G_{\bm{\phi}}(\bm{\phi_{0}},\bm{\theta}_{0})=\begin{bmatrix}0&-\frac{1}{\lambda_{0}^{2}}h_{3}(\lambda_{0})\\[2.84526pt] -2h_{35}(\lambda_{0})&0\end{bmatrix},\,\,I_{\phi,\phi}(\bm{\phi}_{0},\bm{\theta}_{0})=\begin{bmatrix}4(\lambda_{0}+1)~&~0\\[2.84526pt] 0~&~\frac{1}{\lambda_{0}^{3}}\left\{\Big(\frac{1}{\lambda_{0}}+1\Big)\psi_{1}\Big(\frac{1}{\lambda_{0}}+1\Big)-1+(C_{\lambda_{0}}+1)^{2}\right\}\end{bmatrix}$,
$I_{\phi,\theta}(\bm{\phi}_{0},\bm{\theta}_{0})=\begin{bmatrix}\frac{-2\lambda_{0}^{2-1/\lambda_{0}}}{\Gamma(1/\lambda_{0})}&0\\[2.84526pt] 0&-\frac{1}{\lambda_{0}}(C_{\lambda_{0}}+1)\end{bmatrix},\,\,\Sigma_{\mathcal{R}}(\bm{\phi}_{0},\bm{\theta}_{0})=\begin{bmatrix}4(\lambda_{0}+1)-\displaystyle\frac{4\lambda_{0}^{2}}{\Gamma(2-1/\lambda_{0})\,\Gamma(1/\lambda_{0})}&0\\[2.84526pt] 0&\frac{\big(\frac{1}{\lambda_{0}}+1\big)\psi_{1}\big(\tfrac{1}{\lambda_{0}}+1\big)-1}{\lambda_{0}^{3}}\end{bmatrix}$,
$S_{\phi,\theta}(\bm{\phi}_{0},\bm{\theta}_{0})=\begin{bmatrix}\frac{-4\Gamma(2/\lambda_{0})}{\lambda_{0}^{1/\lambda_{0}}\Gamma(3/\lambda_{0})}&0\\[2.84526pt] 0&-\frac{2D_{\lambda_{0}}}{\lambda_{0}^{2}}\left\{2\ln(\lambda_{0})+3\psi(3/\lambda_{0})-\psi(1/\lambda_{0})\right\}\end{bmatrix}$,
where $C_{\lambda_{0}}=\psi(1/\lambda_{0}+1)+\ln(\lambda_{0})$, $\Gamma(\cdot)$, $\psi(\cdot)$ and $\psi_{1}(\cdot)$ denote the gamma, digamma and trigamma functions, respectively, and the functions $h_{i}(\cdot)$ are defined in Table 13. Note that the resulting matrices $V(\bm{\phi}_{0},\bm{\theta}_{0})$ (for both the ML and MM cases) and $\Sigma_{\mathcal{R}}(\bm{\phi}_{0},\bm{\theta}_{0})$ depend only on the value of $\lambda_{0}$. Therefore, to simplify the expressions, we set $\mu_{0}=0,\sigma_{0}=1$.
@@
See Section S14 of the Omnibus goodness-of-fit tests for univariate continuous distributions based on trigonometric moments.
@@
Note that Rao’s score test for the EPD model coincides with the goodness-of-fit test based on the $\lambda$-th power skewness and kurtosis introduced by Desgagné et al. [10]. Since the family of local alternatives is the APD, which accommodates a wide range of asymmetry and tail thickness, this test can be used as an omnibus procedure even though it is specifically designed for this APD alternative. This helps explain why the power curves of the $T_{n}$ tests and of the GLRT/score-test benchmark are much closer for the EPD case than for the gamma case, for which the family of alternatives is more restrictive.
6 Illustrative application to temperature forecast errors